OpenStudy (anonymous):

please check...solve the equations a.2x-30=20 b. (x-8)(x+6)=0 c.x^2+6x=7 check thread for answers.

OpenStudy (anonymous):

a.x=25 b.x=8,-6 c.=-7,1

OpenStudy (anonymous):

i have another one but i cant figure it out..it is 2x^3=-16

OpenStudy (anonymous):

well the responses you got are all correct. As for this new problem, you need to use an algebraic identity (a+b)^3 = a^3 +(3)(a^2)(b)+(3)(a)(b^2) + b^3

OpenStudy (anonymous):

ok so do i start the problem out with those steps?

OpenStudy (anonymous):

i know i have to find x

OpenStudy (anonymous):

would it be -2?